# Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenario

### v1

@@ V Conclusion
In conclusion, we have clarified that the chiral phase transition in QCD-like theories can be Coleman-Weinberg type supercooling. This possibility paves a way to a new search window for new physics to be probed by gravitational wave and primordial black hole productions. We have discussed this possibility by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model in hot and/or dense medium, at the mean-field approximation level. It turned out that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly of the CW-type. The Coleman-Weinberg type chiral phase transition is possible to realize when ordinary QCD is coupled to or replaced by Beyond the Standard Model in a nontrivial way. One candidate scenario includes prediction of an axionlike particle, which is likely to have a rich cosmology and phenomenological consequences at around the QCD scale. More detailed study would be worth performing in the future.
Arguments similar to the present Nambu-Jona-Lasinio’s description can also be applied to bosonic models, e.g., quark meson models and linear sigma models, or more generic boson-fermion system. The key ingredient is at any rate identifying the CW-type scale anomaly and the size of the soft-scale breaking. Such a more general investigation would also be noteworthy.

### v2

@@ V Conclusion
In conclusion, we have clarified that the chiral phase transition in QCD-like theories can be CW type supercooling. This possibility paves a way to a new search window for new physics to be probed by GW and PBH productions. We have discussed this possibility by monitoring ordinary QCD setup in a view of a two-flavor NJL model in hot and/or dense medium, at the MFA level. It turned out that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly of the CW-type. The CW-type chiral-phase transition is possible to realize when ordinary QCD is coupled to or replaced by Beyond the Standard Model in a nontrivial way. A novel scenario includes prediction of an ALP with mass $\gtrsim$ sub GeV, which is likely to have a rich cosmology and phenomenological consequences at around the QCD scale, such as GW and PBH productions. More detailed study would be worth performing in the future.
The CW-type chiral phase transition is realized due to the dominance of the quantum scale anomaly of the logarithmic form at around the chiral criticality. This qualitative feature would still be operative even including gluonic or alternatively non-local (momentum dependent) interactions among quarks, because of persistence of the logarithmic renormalization group running for the potential (more precisely the induced quartic coupling of the CW scalar, which is identified as the bosonized quark bilinear in QCD-like theories). This characteristic logarithmic scale-anomaly is seen in any regularization scheme Kohyama:2015hix , unless dimensional reduction takes place as in the case with a strong magnetic field. Thus, our present proposal would be free from the intrinsic regularization dependence on the quantitative deformation of the chiral phase structure as has been addressed in the literature Kohyama:2015hix .
Arguments similar to the present NJL description can also be applied to bosonic models, e.g., quark meson models and linear sigma models, or more generic boson-fermion system. The key ingredient is at any rate identifying the CW-type scale anomaly and the size of the soft-scale breaking. Such a more general investigation would also be noteworthy.